# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

### v1

@@ Abstract
Most of the approaches proposed so far to craft targeted adversarial examples against Deep Learning classifiers are highly suboptimal and typically rely on increasing the likelihood of the target class, thus implicitly focusing on one-hot encoding settings. In this paper, we propose a more general, theoretically sound, targeted attack that resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction. The minimization is solved by exploiting the Wolfe duality theorem, reducing the problem to the solution of a Non-Negative Least Square (NNLS) problem. The proposed algorithm provides an optimal solution to a linearized version of the adversarial example problem originally introduced by Szegedy et al. [1]. The experiments we carried out confirm the generality of the proposed attack which is proven to be effective under a wide variety of output encoding schemes. Noticeably, the JMA attack is also effective in a multi-label classification scenario, being capable to induce a targeted modification of up to half the labels in a complex multilabel classification scenario with 20 labels, a capability that is out of reach of all the attacks proposed so far. As a further advantage, the JMA attack usually requires very few iterations, thus resulting more efficient than existing methods.
@@ 1 Introduction
Adversarial examples, namely, quasi-imperceptible perturbations capable to induce an incorrect decision, are a serious threat to deep-learning classifiers [1, 2]. Since the publication of the seminal work by Szegedy et al. [1] in which the existence of adversarial examples was first observed, a large number of gradient-based methods have been proposed to implement adversarial attacks against Deep Neural Networks (DNNs) in white-box and black-box scenarios. Most of the attacks proposed so far rely on greedy algorithms that permit to find an effective adversarial example in a reasonable amount of time [2, 3, 4, 5, 6, 7].
Most attacks create the adversarial examples by minimizing a cost function subject to a constraint on the maximum perturbation introduced in the image. Moreover, they focus on classifiers based on one-hot encoding. In this setting, the final activation layer consists in the application of an activation function (usually a softmax) and a normalization, mapping the last layer outputs, called logits, into a probability vector, associating a probability value to each class. The loss function corresponds to the categorical cross entropy. As a consequence, in the targeted case, adversarial examples focus on raising the probability value of the target class. This is obviously the best strategy with one-hot encoding in the presence of a softmax function, however, this strategy is no more optimal in the presence of output encoding schemes based on channel coding, and in the case of multi-label classification.
@@
A more flexible attack working at the logits level has been proposed by Carlini and Wagner [7]. The attack works by decreasing the difference between the largest logit and the logit of the target class (in the targeted case). Working at the logits level allows to avoid vanishing gradient problems [8], hence Carlini and Wagner method often yields better performance compared to methods that directly minimize the loss term. However, working on two logits at a time, without considering the effect of the perturbation on the other logits, is clearly suboptimum in general, with the consequence that the approach in [7] often results in a greedy, lengthy, iterative minimization process. The problem is more evident when the attacker aims at attacking a DNN adopting different encoding mechanisms, like Error Correction Output Coding (ECOC) [9], and also multi-label DNNs. In this case, all the logits have to be modified simultaneously.
@@
Crafting adversarial examples for networks adopting a generic output encoding scheme, and for multi-label classification, is a very challenging task, that, to the best of our knowledge, has not been given much attention, with the exception of some scattered works proposing suboptimum approaches [10, 11]. From the attacker side, the difficulties of applying an optimal attack in this case are due to the fact that the labels are not mutually exclusive (like in the one- hot case), and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way.
@@
To overcome the above drawbacks, we introduce a new, theoretically sound, targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA). JMA solves the original formulation of the adversarial example problem introduced by Szegedy et al. [1], which aims at minimizing the strength of the adversarial perturbation subject to the constraint that the image is classified as belonging to the target class. The solution is found via a two-step procedure: i) given a target point in the output space, the optimum perturbation moving the input sample to the target point is first determined, under a linear assumption on the effect of the distortion on the model output; ii) then, the target point that minimizes the perturbation introduced by the attack is determined by minimizing the Mahalanobis distance induced by the Jacobian matrix of the network input-output function. By exploiting the Wolfe duality theorem, we reduce the problem to the solution of a Non-Negative Least Square (NNLS) problem, that can be efficiently solved numerically. The optimal perturbation derived in this way is then applied to the input image. Ideally, the JMA attack should produce the desired adversarial image in just one-shot. However, due to the local optimality of the proposed approach, in practice, it is sometimes necessary to carry out some iterations to obtain a valid adversarial example, recomputing the Jacobian matrix every time.
@@
The experiments we carried out confirm that the proposed attack is more efficient than state-of-the-art attacks from a computational perspective, requiring a much lower number of iterations to carry out the attack. State-of-the-art attacks are also outperformed in terms of distortion and attack success rate. The effectiveness of the proposed adversarial attack method is particularly evident when it is used to attack networks adopting Error Correction Output Coding (ECOC) [9] and in multi-label classification scenarios, where the JMA attack is capable to simultaneously change in a desired way up to 10 out of 20 labels of the output label vector, a capability which is out of reach of the algorithms proposed so far. We also verified that the method has a very good behavior in the one-hot encoding scenario, where it achieves performance that are comparable to the state-of-the-art with a significant reduction of the computing time. In summary, the main contributions of the paper are:
@@
  • We introduce a new targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA). The attack resorts to the minimization of a Jacobian-induced MAhalanobis distance (JMA) term, with the Jacobian matrix taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction. In theory, the algorithm is one-shot.
@@
  • We validate the new attack considering several datasets (CIFAR-10, GTSRB, MNIST, VOC2012) and networks adopting different encoding schemes, namely the ECOC scheme, multi-label classification, and also one-hot encoding. Our experiments confirm that the proposed attack is very general and can work in all these cases, being very efficient, often requiring only very few iterations to attack the image.
@@
  • We compare JMA against several state-of-the-art methods, showing that the proposed attack is more efficient and requires a much lower number of iterations. In the cases of ECOC and multi-label classification, it also achieves lower distortion and higher attack success rate. In particular, in the multi-label classification scenario, the proposed attack is capable to change up to half labels of the sample output vector, even when the number of labels increases.
@@
The rest of the paper is organized as follows: in Section 2, we provide the main formalism and describe the various classification frameworks addressed in the paper. In Section 3, we review the most relevant adversarial attacks introduced so far, with particular attention to the attacks against DNNs adopting with output channel encoding. Then, in Section 4, we present the JMA adversarial attack. The experimental methodology and setting are described in Section 5. Sections 6 and 7 report and discuss the results of the experiments. The papers ends in Section 8 with our final conclusions.
2 Background and Notation
@@
Let $x\in\mathbb{R}^{m}$ denote the input of the network, and $y(x)$ (or simply $y$) the class $x$ belongs to. We indicate with $l$ the number of classes. To classify $x$, the neural network first maps $x$ into a reduced dimensional space $\mathbb{R}^{n}$ ($n<m$). Every class is associated to an output label column vector $c_{k}=(c_{k1},c_{k2},...,c_{kn})^{T}$. Let $f(x):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ be a column vector indicating the end-to-end neural network function, and $f_{i}(x)$ the $i$-th element of the vector. Classification in favour of one of the $l$ classes is obtained by applying a function $\phi$ to $f(x)$, the exact form of $\phi$ depending on the output encoding scheme used by the network. In the following, we denote by $z=(z_{1},z_{2},...,z_{n})^{T}$ the vector with the logit values, that is, the values of the network nodes before the final activation function, which is responsible to map the output of the penultimate layer of the network into the [0,1] (sometimes [-1, 1]) range. Given an image $x$, the notation $f^{-j}(x)$ is used to refer to the internal model representation at layer $L-j$ ($L$ being the total number of layers of the network). With this notation, $z(x)=f^{-1}(x)$.
2.1 One-hot encoding
@@
Typically, in the one-hot encoding case, training is carried out by minimizing the categorical cross-entropy loss defined as $\mathcal{L}(x,y)=-\sum_{i=1}^{l}c_{yi}\log(f_{i}(x))=-\log f_{y}(x)$ [^1] We remind that $y$ denotes the class sample $x$ belongs to..
2.2 Error-correction-output-coding (ECOC)
@@
Formally, with ECOC, a distinct codeword $c_{k}$ is assigned to every output class. Let $C=\{c_{1},c_{2},\cdots,c_{l}\}$ define the codebook, that is, the matrix of codewords. Each element of $C$ can take values in $\{-1,1\}$. To compute the output of the network the logits are mapped into the $[-1,1]$ range by means of an activation function $\sigma()$, that is $f(x)=\sigma(z)$, where $\sigma()$ is applied element-wise to the logits. A common choice for $\sigma()$ is the $\tanh$ function. To make the final decision, the probability of class $k$ is first computed:
@@
where $f(x)^{T}c_{k}$ is the inner product between $f(x)$ and $c_{k}$. Since $c_{ij}$’s take values in $\{-1,1\}$, the max is necessary to avoid negative probabilities. Then, the model’s final prediction is given by $\phi(x)={\arg\max}_{k}\hskip 2.84544ptp_{k}(x)$. Note that Eq. (2) reduces to Eq. (1) in the case of one-hot encoding, when ${C}=\mathbf{I}_{l}$ and where $\mathbf{I}_{l}$ is the identity $l\times l$ matrix.
In the ECOC case, training is usually carried out by minimizing the hinge loss function, defined as $\mathcal{L}(x,y)=\sum_{i=1}^{n}\max(1-c_{yi}f_{i}(x),0)$.
2.3 Multi-label classification
@@
In this paper, we refer to multi-label classification as a scenario wherein the classifier is asked to decide about the presence of absence within the image of $n$ image characteristics or features. For example, the classifier may be asked to detect the presence of $n$ possible classes of objects, or decide about $n$ binary attributes (like indoor/outdoor, night/day, sunny/rainy). The presence/absence of the looked-for features is encoded by the $n$ outputs of the network (see, for instance, [13, 14]). In general, the number of possible outputs is $l=2^{n}$, and each output is encoded in a matrix $C$ having size $n\times 2^{n}$. In fact, it is possible that some labels’ combinations are not feasible. In such a case, $C$ contains only the allowed combinations, somewhat playing the same role of channel coding. In this paper, by multi-label classification, we indicate a situation where all combinations of labels are possible.
@@
In the case of multi-label classification, given the logit vector $z$, the activation function $\sigma()$ is applied element-wise to the components of $z$, and the prediction on each label component is made component-wise. Assuming that $c_{k}\in\{0,1\}^{n}$ for any $k$, and that the logistic function is applied to the logits, decoding corresponds to 0.5-thresholding each output score independently [^2] If $c_{k}\in\{-1,1\}^{n}$, the tanh activation is applied followed by 0-thresholding..
For the loss function, a common choice is the binary cross-entropy loss $\mathcal{L}(x,y)=-\sum_{i=1}^{n}(c_{yi}\log(f_{i}(x))+(1-c_{yi})\log(1-f_{i}(x)))$.
3 Adversarial attacks against DNNs

### v2

@@ Abstract
Most of the approaches proposed so far to craft targeted adversarial examples against Deep Learning classifiers are highly suboptimal and typically rely on increasing the likelihood of the target class, thus implicitly focusing on one-hot encoding settings. In this paper, a more general, theoretically sound, targeted attack is proposed, which resorts to the minimization of a Jacobian-induced Mahalanobis distance term, taking into account the effort (in the input space) required to move the latent space representation of the input sample in a given direction. The minimization is solved by exploiting the Wolfe duality theorem, reducing the problem to the solution of a Non-Negative Least Square (NNLS) problem. The proposed algorithm (referred to as JMA) provides an optimal solution to a linearised version of the adversarial example problem originally introduced by Szegedy et al. The results of the experiments confirm the generality of the proposed attack which is proven to be effective under a wide variety of output encoding schemes. Noticeably, JMA is also effective in a multi-label classification scenario, being capable to induce a targeted modification of up to half the labels in complex multi-label classification scenarios, a capability that is out of reach of all the attacks proposed so far. As a further advantage, JMA requires very few iterations, thus resulting more efficient than existing methods.
@@ I Introduction
Adversarial examples, namely, quasi-imperceptible perturbations capable to induce an incorrect decision, are a serious threat to deep-learning classifiers [1, 2]. Since the publication of the seminal work by Szegedy et al. [1] in which the existence of adversarial examples was first observed, a large number of gradient-based methods have been proposed to implement adversarial attacks against Deep Neural Networks (DNNs) in white-box and black-box scenarios [2, 3, 4, 5].
@@
Most attacks create the adversarial examples by minimizing a cost function subject to a constraint on the maximum perturbation introduced in the image. Moreover, they focus on single-label classifiers based on one-hot encoding. In this setting, the final activation layer consists in the application of an activation function (usually a softmax) and a normalization, mapping the last layer outputs, called logits, into a probability vector, associating a probability value to each class. The loss function corresponds to the categorical cross entropy. As a consequence, in the targeted case, adversarial examples focus on raising the probability value of the target class. This is obviously the best strategy with single-label classification however, this strategy is not optimal in general, e.g., in the presence of output encoding schemes based on channel coding, and in the case of multi-label classification. A more flexible attack working at the logits level has been proposed by Carlini and Wagner [6]. The attack works by decreasing the difference between the largest logit and the logit of the target class (in the targeted case). Working at the logits level allows to avoid vanishing gradient problems [7], hence Carlini and Wagner method often yields better performance compared to methods that directly minimize the loss term. However, working on two logits at a time, without considering the effect of the perturbation on the other logits, is clearly suboptimum in general, with the consequence that the approach in [6] often results in a greedy, lengthy, iterative minimization process. The problem is more evident when the attacker aims at attacking a DNN adopting different encoding mechanisms, like Error Correction Output Coding (ECOC) [8], and also multi-label DNNs. In this case, all the logits have to be modified simultaneously.
@@
Crafting adversarial examples for networks adopting a generic output encoding scheme, and for multi-label classification, is a very challenging task, that, to the best of the authors’ knowledge, has not been given much attention, with the exception of some scattered works proposing suboptimum approaches [9, 10], [11]. From the attacker side, the difficulties of applying an optimal attack in this case are due to the fact that the labels are not mutually exclusive (like in the one- hot case), and decreasing or increasing the network output in correspondence of some nodes may change the values taken by the other nodes in an unpredictable way.
@@
To overcome the above drawbacks, we introduce a new, theoretically sound, targeted adversarial attack, named Jacobian-induced Mahalanobis distance Attack (JMA). JMA solves the original formulation of the adversarial example problem introduced in [1], which aims at minimizing the strength of the adversarial perturbation subject to the constraint that the image is classified as belonging to the target class. The solution is found via a two-step procedure: i) given a target point in the output space, the optimum perturbation moving the input sample to the target point is first determined, under a linear assumption on the effect of the distortion on the model output; ii) then, the target point that minimizes the perturbation introduced by the attack is determined by minimizing the Mahalanobis distance induced by the Jacobian matrix of the network input-output function. By exploiting the Wolfe duality theorem, the problem is reduced to the solution of a Non-Negative Least Square (NNLS) problem, that can be efficiently solved numerically. The optimal perturbation derived in this way is then applied to the input image. Ideally, JMA should produce the adversarial image in just one-shot. However, due to local optimality, in practice, it is sometimes necessary to carry out some iterations to obtain a valid adversarial example, recomputing the Jacobian matrix every time.
@@
The results of the experiments confirm that the proposed attack is more efficient than state-of-the-art attacks from a computational perspective, requiring lower iterations to carry out the attack. State-of-the-art attacks are also outperformed in terms of distortion and attack success rate. The effectiveness of the proposed adversarial attack method is particularly evident when it is used to attack networks adopting ECOC and in particular in multi-label classification scenarios, where JMA is capable to simultaneously change in a desired way up to 10 out of 20 labels of the output label vector, a capability which is out of reach of the algorithms proposed so far. We also verified the very good behavior of JMA in the one-hot encoding scenario, where it achieves performance that are comparable to the state-of-the-art with a significant reduction of the computing time. In summary, the main contributions of this paper are:
@@
  • We propose a new targeted adversarial attack for general classification frameworks, named Jacobian-induced Mahalanobis distance Attack (JMA). The attack resorts to the minimization of a Jacobian-induced Mahalanobis distance term, with the Jacobian matrix taking into account the effort (input space) to move the latent space representation of the input sample in a given direction. In theory, the algorithm is one-shot.
@@
  • We validate the new attack on various datasets (CIFAR-10, GTSRB, MNIST, VOC2012, MS-COCO, NUS-WIDE, and ImageNet) and networks adopting different encoding schemes, namely ECOC, multi-label, and one-hot encoding. Validation involves different types of architectures, including CNNs, Transformers and large visual models. The experiments confirm that the proposed attack is very general and can work in all these cases, being very efficient, often requiring only very few iterations to attack an image.
@@
  • We compare our new approach against several state-of-the-art attacks, showing that JMA is more efficient and requires less iterations. In particular, in the multi-label classification scenario, JMA is capable to change up to half labels of the output vector, even when the number of labels increases.
@@
The rest of the paper is organized as follows: Section II gives an overview of the related work. Section III provides the main formalism and introduces the various classification frameworks. In Section IV, the most relevant adversarial attacks that can be applied against DNNs adopting output encoding are reviewed. Then, in Section V, the details of the JMA adversarial attack are described. The experimental methodology and setting are described in Section VI, while section VII reports and discusses the results of the experiments. The papers ends in Section VIII with some final remarks.
II Related works
Adversarial attacks can be categorized in two main groups, namely white-box and black-box methods [5].
Starting from Szegedy et al. [1] work, research on white-box attacks has mainly focused on the development of gradient-based approaches that can reduce the complexity of the attack. Many greedy algorithms have been proposed that permit to find an effective adversarial example in a reasonable amount of time [2, 3, 12, 13, 14, 6]. The Fast Sign Gradient Method (FGSM) method [2] obtains an adversarial perturbation in a computationally efficient way by considering the sign of the gradient of the output with respect to the input image. In [12], an iterative version of FGSM is introduced by applying FGSM multiple times, with a smaller step size, each time by recomputing the gradient. This method is often referred to as I-FGSM, or Basic Iterative Method (BIM). Another attack similar to FGSM is the projected gradient descent (PGD) attack [13], that can be regarded to as a multi-step extension of the FGSM attack where the clip operation performed by BIM on the gradient (to force the solution to stay in the [0,1] range) is replaced by gradient projection. Carlini and Wagner (C&W) [6] propose a more flexible method, working at the logits level, which can mitigate the gradient vanishing problem and improve performance in many cases. Recently, the AutoPGD and AutoAttack [15] attacks have also been proposed. Such attacks improve PGD by designing an approach that automatically chooses the most suitable step size and perturbation size (in the case of AutoAttack) at every attack iteration, based on the behaviour of the objective function, and by using a different loss function, which improves performance and reduces the gradient vanishing problem. Among the other white-box gradient-based attacks it is worth mentioning the Jacobian-based Saliency Map Attack (JSMA) [3] and the DeepFool attack [14]. JSMA [3] consists of a greedy iterative procedure which relies on forward propagation to compute, at each iteration, a saliency map, indicating the pixels that contribute most to the classification, while DeepFool [14] is an efficient iterative attack that considers the minimal perturbation with respect to a linearized classifier, stopping the attack when the boundary is crossed.
A large segment of recent literature has also focused on the development of attacks that can work in real-world settings where the exact victim model is unknown. This includes transfer-based attacks, e.g. [4, 16, 17], which retain part of their effectiveness even against DNN models other than the one targeted by the attack, as well as black-box attacks operating under the assumption that the victim model can be queried a limited number of times [18, 19].
All the above methods focus on single-label classifiers adopting one-hot encoding schemes. The problem of crafting adversarial attacks against classifiers adopting different output encoding schemes is a challenging one that, to the best of the authors’ knowledge, has not been much studied, with the exception of a few scattered works. Song et. al [9] first addressed this problem and extended the C&W and the DeepFool attacks to a multi-label setting. These extended algorithms are referred as ML-C&W and ML-DF. Inspired by ML-DF, another approach to implement an adversarial attack in the multi-label case, named Multi-Label Attack via Linear programming (MLA-LP), has been proposed in [10]. Recently, [11] proposed a fundamentally different multi-label attack method that enforces semantic consistency across all predicted labels in the adversarial image. The approach introduced in [11] leverages an efficient search algorithm over a knowledge graph that captures label dependencies. Beyond multi-label classification, a white-box attack tailored against DNNs adopting the ECOC encoding scheme has been proposed in [20]. A completely different approach, that can be applied to any network regardless of the encoding scheme, often leading to a large attack distortion in the input space, is the Layerwise Origin-Target Synthesis (LOTS) attack [21]. Due to their relevance for our research, the attacks targeting classifiers adopting output encoding are presented in detail in Section IV-B.
III Background and Notation
@@
Let $x\in\mathbb{R}^{m}$ denote the input of the network, and $y(x)$ (or simply $y$) the class $x$ belongs to. Le $l$ be the number of classes. To classify $x$, the neural network first maps $x$ into a reduced dimensional space $\mathbb{R}^{n}$ ($n<m$). Every class is associated to an output label column vector $c_{k}=(c_{k1},c_{k2},...,c_{kn})^{T}$. Let $f(x):\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ be a column vector indicating the end-to-end neural network function, and $f_{i}(x)$ the $i$-th element of the vector. Classification in favour of one of the $l$ classes is obtained by applying a function $\phi$ to $f(x)$, the exact form of $\phi$ depending on the output encoding scheme used by the network. In the following, $z=(z_{1},z_{2},...,z_{n})^{T}$ denotes the vector with the logit values, that is, the values of the network nodes before the final activation function, which is responsible to map the output of the penultimate layer of the network into the [0,1] (sometimes [-1, 1]) range. Given an image $x$, the notation $f^{-j}(x)$ is used to refer to the internal model representation at layer LjL-j ($L$ being the total number of layers of the network). With this notation, $z(x)=f^{-1}(x)$.
III-A One-hot encoding
@@
Typically, in the one-hot encoding case, training is carried out by minimizing the categorical cross-entropy loss defined as $\mathcal{L}(x,y)=-\sum_{i=1}^{l}c_{yi}\log(f_{i}(x))=-\log(f_{y}(x))$.
III-B Error-correction-output-coding (ECOC)
@@
Formally, with ECOC, a distinct codeword $c_{k}$ is assigned to every output class. Let $C=\{c_{1},c_{2},\cdots,c_{l}\}$ define the codebook, that is, the matrix of codewords. Each element of $C$ can take values in $\{-1,1\}$. To compute the output of the network the logits are mapped into the [1,1][-1,1] range by means of an activation function $\sigma()$, that is $f(x)=\sigma(z)$, where $\sigma()$ is applied element-wise to the logits. A common choice for $\sigma()$ is the $\tanh$ function. To make the final decision, the probability of class $k$ is first computed:
@@
where $f(x)^{T}c_{k}$ is the inner product between $f(x)$ and $c_{k}$. Since $c_{ij}$’s take values in $\{-1,1\}$, the max is necessary to avoid negative probabilities. Then, the model’s final prediction is given by $\phi(x)={\arg\max}_{k}\hskip 2.84544ptp_{k}(x)$. Note that Eq. (2) reduces to Eq. (1) in the case of one-hot encoding, when ${C}=\mathbf{I}_{l}$ and where $\mathbf{I}_{l}$ is the identity $l\times l$ matrix. In the ECOC case, training is usually carried out by minimizing the hinge loss function, defined as $\mathcal{L}(x,y)=\sum_{i=1}^{n}\max(1-c_{yi}f_{i}(x),0)$.
III-C Multi-label classification
@@
Multi-label classification is the scenario wherein the classifier is asked to decide about the presence or absence within the image of $n$ image characteristics or features. For example, the classifier may be asked to detect the presence of $n$ possible classes of objects, or decide about $n$ binary attributes (like indoor/outdoor, night/day, sunny/rainy). The presence/absence of the looked-for features is encoded by the $n$ outputs of the network (see, for instance, [23, 24]). In general, the number of possible outputs is $l=2^{n}$, and each output is encoded in a matrix $C$ having size $n\times 2^{n}$. In fact, it is possible that some labels’ combinations are not feasible. In such a case, $C$ contains only the allowed combinations, somewhat playing the same role of channel coding. In this paper, multi-label classification indicates a situation where all combinations of labels are possible.
@@
In the case of multi-label classification, given the logit vector $z$, the activation function $\sigma()$ is applied element-wise to the components of $z$, and the prediction on each label component is made component-wise. Assuming that $c_{k}\in\{0,1\}^{n}$ for any $k$, and that the logistic function is applied to the logits, decoding corresponds to 0.5-thresholding each output score independently [^1] If $c_{k}\in\{-1,1\}^{n}$, tanh activation followed by 0-thresholding is applied.. For the loss function, a common choice is the (multi-label) binary cross-entropy loss $\mathcal{L}(x,y)=-\sum_{i=1}^{n}(c_{yi}\log(f_{i}(x))+(1-c_{yi})\log(1-f_{i}(x)))$.
IV Adversarial attacks against DNNs